Fermi National Laboratory
 

Operation Summary for March 16, 2011 - March 18, 2011

 

Wednesday March 16

The day shift began with Operations monitoring the Tevatron (TEV) store 8587 (156.43E30) and the Recycler stash (82.6E10), with the Antiproton Source (Pbar) stacking (10.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Booster personnel replaced a corrector magnet power supply regulator (HL11).

During the evening shift, at 7 PM, a Recycler kicker (KPS2A) tripped off on a trigger ready indication. The Run Coordinator decided to terminate store 8587 early so as to not risk losing the 314E10 antiprotons in the Recycler due to a KPS2A pre-fire. Operations terminated store 8587 at 7:35 PM, put the TEV into HEP shot setup, and established store 8588 at 8:50 PM with an average initial luminosity (AIL) of 331E30. An EE Support technician investigated the KPS2A problem after the store had been established and adjusted the high voltage setting on a control board. Operators filled the Pbar dump water system once during the shift.

Thursday March 17

The midnight shift began with Operations monitoring store 8588 (174.98E30) and the stash (110.7E10), with Pbar stacking (29.5E10), and with MiniBooNE taking beam.

During the midnight shift, operators filled the Pbar dump water system. Operations reported that a Linac RF station (LRF2) tripped off several times due to an increase in the power amplifier vacuum pressure.

During the day shift, Operations terminated store 8588 at noon, put the TEV into shot setup, and established store 8589 at 1:10 PM with an AIL of 367E30. FESS personnel reported that they had configured the MI-30 LCW pump to run in remote. Operations reported that the Recycler kicker KPS2A tripped off just before shot setup with the same fault as seen on the previous evening shift. Kicker experts started working on KPS2A after the store was in the TEV.

During the evening shift, Operators added water to the Pbar dump water system.

Friday March 18

The midnight shift began with Operations monitoring store 8589 (105.51E30) and the stash (268.8E10), with Pbar stacking (8.2E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8589 at 6:35 AM, put the TEV into shot setup, and established store 8590 with an AIL of 359E30.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/18/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory